Black iron (III) oxide solid can be converted into water and iron metal when it is reacted with hydrogen gas. If 125 g of iron (III) oxide reacted, how many moles of iron can be obtained?
1
Expert's answer
2020-10-20T08:37:26-0400

Fe2O3 + 3H2 → 2Fe + 3H2O

M(Fe2O3) = 159.69 g/mol

n = m/M

n(Fe2O3) "= \frac{125}{159.69}=0.783\; mol"

n(Fe) = 2n(Fe2O3) "= 2\times 0.783 = 1.566 \;mol"

Answer: 1.56 mol